# 3D-LLaVA: Towards Generalist 3D LMMs with Omni Superpoint Transformer

### v1

@@ Abstract
Current 3D Large Multimodal Models (3D LMMs) have shown tremendous potential in 3D-vision-based dialogue and reasoning. However, how to further enhance 3D LMMs to achieve fine-grained scene understanding and facilitate flexible human-agent interaction remains a challenging problem. In this work, we introduce 3D-LLaVA, a simple yet highly powerful 3D LMM designed to act as an intelligent assistant in comprehending, reasoning, and interacting with the 3D world. Unlike existing top-performing methods that rely on complicated pipelines—such as offline multi-view feature extraction or additional task-specific heads—3D-LLaVA adopts a minimalist design with integrated architecture and only takes point clouds as input. At the core of 3D-LLaVA is a new Omni Superpoint Transformer (OST), which integrates three functionalities: (1) a visual feature selector that converts and selects visual tokens, (2) a visual prompt encoder that embeds interactive visual prompts into the visual token space, and (3) a referring mask decoder that produces 3D masks based on text description. This versatile OST is empowered by the hybrid pretraining to obtain perception priors and leveraged as the visual connector that bridges the 3D data to the LLM. After performing unified instruction tuning, our 3D-LLaVA reports impressive results on various benchmarks. The code and model will be released to promote future exploration.
@@ 1 Introduction
Figure 1: An intuitive comparison between 3D-LLaVA and other SoTA 3D LMMs. 3D-LLaVA achieves the best performance among the competitors on most of the benchmarks.
@@
To overcome these limitations, we present 3D-LLaVA, a generalist 3D LMM that streamlines the pipeline while maintaining strong performance across diverse 3D tasks. In contrast to the prior works that assemble multiple models or extract features in an offline manner, 3D LLaVA bridges interactive 3D vision dialogue and point-level 3D scene comprehension in an integrated and shared architecture, eliminating the need for auxiliary modules and complicated steps. Particularly, as compared in Figure 1, the most distinguishing part of 3D-LLaVA is the novel visual connector, namely Omni Superpoint Transformer (OST). What makes it distinguishing includes how we use it as a shared module for multiple purposes and how we pretrain it together with the 3D scene encoder.
@@
Moreover, at the pertaining stage, OST is connected together with the 3D scene encoder and jointly pre-trained with the hybrid supervision of instance segmentation and 2D-to-3D knowledge distillation. Here, the 2D feature is extracted from multi-view images with the visual encoder of a 2D LMM, i.e. LLaVA-1.5 [42], and lifted to 3D by the geometric correspondence [46] between the point cloud and the pixels. Such a pretraining scheme on the one hand encompresses the perception prior to our model, and takes the well-aligned 2D data as the bridge to facilitate the alignment between 3D visual embedding and language embedding.
@@ 3 Approach
The overall framework of 3D-LLaVA is illustrated in Figure 2. It is a generalist 3D LMM, capable of conducting 3D vision-centric dialogue, being interacting seamlessly with flexible visual and textual prompts, and grounding open-ended language description into 3D point cloud masks. In this section, we first introduce the model architecture of the 3D scene encoder (Section 3.1) and Omni Superpooint Transformer (Section 3.2). Then, in Section 3.3, we elaborate on the details of each step in our pipeline. Finally, in Section 3.4, we introduce the training scheme.
@@ 3.2 Omni Superpoint Transformer
There are three heads on the top of OST: a mask head, a classification head, and an alignment head. The mask head transforms each query embedding into a mask prediction kernel, which is then applied to generation binary mask prediction by performing a dot product with the input superpoint features of OST [53, 52, 29]. The classification head predicts the category of the segmentation mask by outputting the logit of each category. The output of the alignment head is denoted as $Z_{V}$ in Figure 2. It would be further leveraged to obtain the visual token of the LLM.
@@ 3.3 Details in Pipeline
Visual Prompt Encoding. A generalist 3D LMM is supposed to be interacted with by both language instructions and visual prompts. Common visual prompts include a clicking point, bounding box, or binary mask. A straightforward approach to encode these prompts is to use a prompt encoder composed of several linear layers [11], designed to project the prompt (e.g., coordinates of clicking points or bounding boxes) into an embedding that aligns with the same semantic space as visual or language tokens. However, we find that this type of prompt encoder is challenging to optimize, as it lacks explicit information indicating which areas are targeted by the prompt.
@@
Instruction Formulation. We present the typical language instruction in the bottom-right part of Figure 2. There are two kinds of place holders in the instruction: “⟨PC⟩” and “⟨Visual Prompt⟩”. The text instruction except for the placeholder will be tokenized into the text token embedding $H_{T}$. After tokenization, we replace “⟨PC⟩” with visual token embedding $H_{V}$, and replace “⟨Visual Prompt⟩” with the prompt token embedding $H_{P}$.
@@ 3.4 Training Scheme
Stage 1: Pre-training 3D Scene Encoder and OST.
Unlike the 2D domain, which has powerful and widely recognized vision foundation models such as CLIP [50], there is currently no such 3D foundation model that can serve as a readily usable 3D visual encoder.
@@
Others captioning, object description 32K
Total - 393K
Stage 2: End-to-End Instruction Tuning.
We combine various 3D vision and language understanding datasets to form our instruction-tuning data. The combined datasets include ScanRefer [7], ReferIt3D (i.e., Nr3D and Sr3D) [1], ScanQA [3], SQA3D [44], Scan2Cap [13], Multi3DRefer [63], ScanNet scene-level captioning subset of SceneVerse [30] and Object Description subset of Grounded 3D-LLM [12]. The statistic of the utilized dataset is presented in Table 1. To enrich the language annotation that involves describing the object in the 3D scene, we follow [24] to use Nr3D as the complementary to the dense captioning task, which is denoted as “Nr3D*” in this table.
@@
where $L_{text}$ is the cross-entropy loss for next-token generation, $L_{mask}$ represents the mask prediction loss that also consists of the binary cross-entropy loss and the Dice loss, which is the same as the pertaining stage. Here we multiply a coefficient of 0.1 to the mask loss for scale balance. We always keep the Sparse 3D U-Net, OST, and the main body of LLM frozen. Only the visual projector, the SEG project, and LoRA [23] parameters adopted to the LLM are updated.
@@ 4.1 Datasets and Metrics
Datasets. In this work, we conduct experiments on the 3D scans provided by ScanNet dataset [17], including 1,201 scenes for training and 312 for validation. At the pertaining stage for our 3D encoder, we leverage the mask annotation from ScanNet200 [51], which extends the original ScanNet with fine-grained categories. The language annotation leveraged in the instruction tuning has been introduced in Section 4. After instruction tuning, we validate the effectiveness of the proposed 3D-LLaVA on the following datasets: ScanQA [3] and SQA3D [44] for question answering, ScanRefer [7] and Multi3DRefer [63] for referring segmentation and Scan2Cap [13] for dense captioning.
@@ 4.3 Comparison with SoTA Models
3D Referring Segmentation requires the model to output the 3D mask on the point cloud according to the users’ language expression, which validates the capability of grounding the text description on the 3D scene. We benchmark our methods and other state-of-the-art methods on both the single-target setting (Scanerfer [7]) and the various number setting (Multi3DRefer [63]).The referring text from Multi3DRefer can correspond to one, many, or even zero objects. If multiple objects are referred to in the instruction, we follow [20] to merge the masks into a single one for evaluation. When there is no corresponding to the referring expression, our 3D-LLaVA will output “Sorry, I cannot find this object.”. In this case, since there is no [SEG] token in the response, the mask decoding pipeline will not be applied, and thus all of the predicted masks will be assigned as background. As shown in the table, our 3D-LLaVA reports the best result among the competitors. Notably, our model achieves 43.3% mIoU on ScanRefer and 42.7% mIoU on Multi3DRefer, improving the previous best record of SegPoint by absolutely 1.6% mIoU and 6.6% mIoU.
@@
Table 4: Ablation study on the method to select visual tokens. The models are compared in terms of CiDEr and BLEU-4 on ScanQA [3] and Scan2Cap [13]. Our default setting is highlighted with light blue.
ScanQA Scan2Cap
[…]
Random 91.4 16.5 76.8 36.7
Farthest Point Sampling 91.7 16.6 77.7 36.9
Based on OST 92.6 17.1 78.8 36.9
4.4 Ablation Study
In this section, we conduct ablative experiments to analyze the effects of each component in our proposed 3D-LLaVA framework. Unless specified, the model evaluated in this section is trained with the same data and training scheme as the default setting introduced in the former sections.
@@ 4.4 Ablation Study
Effect of Visual Token Selection. In this experiment, we study the effect of the paradigm on selecting visual tokens. The results are shown in Table 4. The most naive choice when selecting visual tokens is to perform random selection. Here we randomly select 100 tokens as the visual input of LLM. The performance drops to 91.4% CiDEr on ScanQA and 76.8% CiDEr on Scan2Cap. Another choice is to apply furthest point sampling (FPS) [47], which is a common operation for point cloud recognition. Compared to random sampling, FPS works better in covering different locations of the 3D scene, improving the CiDER to 91.7% and 77.7% on ScanQA and Scan2Cap, respectively. In our Default setting, we leverage the classification score as the indicator to choose superpoints that are more likely to be the foreground ones, thus guaranteeing the visual token includes the informative region of the whole area.
@@
Figure 5: Visualization of 3D-LLaVA’s response on various tasks. Each of these examples includes instructions on how to perform referring segmentation. Besides, the examples present the result of 3D question answering [3], 3D dense captioning [64], and situated question answering [44], respectively. When the referred object is not in the given 3D scene, the model is aware of responding with “Sorry, I cannot find this object”.

### v2

@@ Abstract
Current 3D Large Multimodal Models (3D LMMs) have shown tremendous potential in 3D-vision-based dialogue and reasoning. However, how to further enhance 3D LMMs to achieve fine-grained scene understanding and facilitate flexible human-agent interaction remains a challenging problem. In this work, we introduce 3D-LLaVA, a simple yet highly powerful 3D LMM designed to act as an intelligent assistant in comprehending, reasoning, and interacting with the 3D world. Unlike existing top-performing methods that rely on complicated pipelines—such as offline multi-view feature extraction or additional task-specific heads—3D-LLaVA adopts a minimalist design with integrated architecture and only takes point clouds as input. At the core of 3D-LLaVA is a new Omni Superpoint Transformer (OST), which integrates three functionalities: (1) a visual feature selector that converts and selects visual tokens, (2) a visual prompt encoder that embeds interactive visual prompts into the visual token space, and (3) a referring mask decoder that produces 3D masks based on text description. This versatile OST is empowered by the hybrid pretraining to obtain perception priors and leveraged as the visual connector that bridges the 3D data to the LLM. After performing unified instruction tuning, our 3D-LLaVA reports impressive results on various benchmarks.
@@ 1 Introduction
Figure 1: An intuitive comparison between 3D-LLaVA and other SoTA 3D LMMs (The performance of LEO on ScanQA is omitted here since its setting is different). Our 3D-LLaVA achieves the best results among the competitors on most of the benchmarks.
@@
To overcome these limitations, we present 3D-LLaVA, a generalist 3D LMM that streamlines the pipeline while maintaining strong performance across diverse 3D tasks. In contrast to the prior works that assemble multiple models or extract features in an offline manner, 3D LLaVA bridges interactive 3D vision dialogue and point-level 3D scene comprehension in an integrated and shared architecture, eliminating the need for auxiliary modules and complicated steps. Particularly, as compared in Figure 1, the most distinguishing part of 3D-LLaVA is the novel visual connector, namely Omni Superpoint Transformer (OST). What distinguishes it is how we use it as a shared module for multiple purposes and how we pretrain it with the 3D scene encoder.
@@
Moreover, at the pretaining stage, OST is connected together with the 3D scene encoder and jointly pre-trained with the hybrid supervision of instance segmentation and 2D-to-3D knowledge distillation. Here, the 2D feature is extracted from multi-view images with the visual encoder of a 2D LMM, i.e. LLaVA-1.5 [42], and lifted to 3D by the geometric correspondence [46] between the point cloud and the pixels. Such a pretraining scheme on the one hand encompasses the perception prior to our model and takes the well-aligned 2D data as the bridge to facilitate the alignment between 3D visual embedding and language embedding.
@@ 3 Approach
The overall framework of 3D-LLaVA is illustrated in Figure 2. It is a generalist 3D LMM, capable of conducting 3D vision-centric dialogue, being interacting seamlessly with flexible visual and textual prompts, and grounding open-ended language description into 3D point cloud masks. In this section, we first introduce the model architecture of the 3D scene encoder (Section 3.1) and Omni Superpooint Transformer (Section 3.2). Then, in Section 3.3, we elaborate on the detail of each step in our pipeline. Finally, in Section 3.4, we introduce the training scheme.
@@ 3.2 Omni Superpoint Transformer
There are three heads on the top of OST: a mask head, a classification head, and an alignment head. The mask head transforms each query embedding into a mask prediction kernel, which is then applied to generate binary mask prediction by performing a dot product with the input superpoint features of OST [53, 52, 30]. The classification head predicts the category of the segmentation mask by outputting the logit of each category. The output of the alignment head is denoted as $Z_{V}$ in Figure 2. It would be further leveraged to obtain the visual token of the LLM.
@@ 3.3 Details in Pipeline
Visual Prompt Encoding. A generalist 3D LMM is supposed to be interacted with both language instructions and visual prompts. Common visual prompts include a clicking point, bounding box, or binary mask. A straightforward approach to encode these prompts is to use a prompt encoder composed of several linear layers [11], designed to project the prompt (e.g., coordinates of clicking points or bounding boxes) into an embedding that aligns with the same semantic space as visual or language tokens. However, we find that this type of prompt encoder is challenging to optimize, as it lacks explicit information indicating which areas are targeted by the prompt.
@@
[…]
@@ 3.4 Training Scheme
Stage 1: Pre-training 3D Scene Encoder and OST. Unlike the 2D domain, which has powerful and widely recognized vision foundation models such as CLIP [50], there is currently no such 3D foundation model that can serve as a readily usable 3D visual encoder.
@@
Total - 295K
Stage 2: End-to-End Instruction Tuning. We combine various 3D vision and language understanding datasets to form our instruction-tuning data. The combined datasets include ScanRefer [7], Nr3D [1], Multi3DRefer [64], ScanQA [3], SQA3D [44], Scan2Cap [13]. The statistic of the utilized dataset is presented in Table 1. To enrich the language annotation that involves describing the object in the 3D scene, we follow [25] to use Nr3D as the complementary to the dense captioning task, which is denoted as “Nr3D*” in this table.
@@
where $L_{text}$ is the cross-entropy loss for next-token generation, $L_{mask}$ represents the mask prediction loss that also consists of the binary cross-entropy loss and the Dice loss, which is the same as the pertaining stage. Here, we multiply the mask loss by a coefficient of 0.1 for balance. We always keep the Sparse 3D U-Net, OST, and the main body of LLM frozen. Only the visual projector, the SEG project, and LoRA [24] parameters adopted to the LLM are updated.
@@ 4.1 Datasets and Metrics
Datasets. In this work, we conduct experiments on the 3D scans provided by ScanNet dataset [17], including 1,201 scenes for training and 312 for validation. At the pertaining stage of our 3D encoder, we leverage the mask annotation from ScanNet200 [51], which extends the original ScanNet with fine-grained categories. The language annotation leveraged in the instruction tuning has been introduced in Section 4. After instruction tuning, we validate the effectiveness of the proposed 3D-LLaVA on the following datasets: ScanQA [3] and SQA3D [44] for question answering, ScanRefer [7] and Multi3DRefer [64] for referring segmentation and Scan2Cap [13] for dense captioning.
@@ 4.3 Comparison with SoTA Models
3D Referring Segmentation requires the model to output the 3D mask on the point cloud according to the user’s language expression, which validates the capability of grounding the text description on the 3D scene. We benchmark our methods and other state-of-the-art methods on both the single-target setting (Scanerfer [7]) and the various number setting (Multi3DRefer [64]).The referring text from Multi3DRefer can correspond to one, many, or even zero objects. If multiple objects are referred to in the instruction, we follow [21] to merge the masks into a single one for evaluation. When there is no corresponding to the referring expression, our 3D-LLaVA will output “Sorry, I cannot find this object.”. In this case, since there is no [SEG] token in the response, the mask decoding pipeline will not be applied, and thus all of the predicted masks will be assigned as background. As shown in the table, our 3D-LLaVA reports the best result among the competitors. Notably, our model achieves 43.3% mIoU on ScanRefer and 42.7% mIoU on Multi3DRefer, improving the previous best record of SegPoint by 1.6% mIoU and 6.6% mIoU.
@@
Table 3: Performance comparisons for box-level 3D visual grounding. Our results are obtained by directly converting the foreground referring mask to a box, highlighted with light blue.
ScanRefer (Box-Level)
Methods Acc@0.25 Acc@0.5
Specialist Models:
3D-VisTA [67] 50.6 45.8
ConcretNet [56] 50.6 46.5
3D LMMs:
3D-LLM [23] 30.3 -
Grounded 3D-LLM [12] 48.6 44.0
Chat-Scene [25] 55.5 50.2
3D-LLaVA (Ours) 51.2 40.6
@@
4.4 Experimental Analysis
This section presents the experimental analysis of our 3D-LLaVA. Unless specified, the model evaluated in this section is trained with the same data and training scheme as the default setting introduced in the former sections.
Evaluating Masks with the Box-level Metric. Even not designed for 3D referring segmentation, our 3D-LLaVA can also produce box-level grounding results. Specifically, we first apply DBSCAN algorithm [19] to the foreground mask to remove outliers, and then obtain the grounding box by considering the minimum and maximum coordinates of the mask. We compare the box-level grounding performance of our 3D-LLaVA with bot the specialist model and other 3D LMMs in Table 3. Although our model is optimized for precise binary masks, whereas competitors are trained to select best-matching proposals based on box IoU, our 3D-LLaVA achieves 51.2% accuracy when the IoU threshold is 0.25, better than most of the competitors in the table. Our performance lags behind Chat-Scene [25], but our method relies on neither an extra mask proposal generator nor the fusion of image and point cloud features.
@@ 4.4 Experimental Analysis
Figure 5: Visualization of 3D-LLaVA’s response on various tasks. Each of these examples includes an instruction to perform referring segmentation. Besides, the examples present the result of 3D question answering [3], 3D dense captioning [65], and situated question answering [44], respectively. When the referred object is not in the given 3D scene, the model is aware of responding with “Sorry, I cannot find this object”.
@@ 5 Conclusion
Acknowledgements: This work was supported by the Centre for Augmented Reasoning, an initiative by the Department of Education, Australian Government.